find the measure of an angle such that the sum of the measures of its complement and suplement is 110
Answer by nerdybill(7384)

find the measure of an angle such that the sum of the measures of its complement and suplement is 110
.
Let x = the angle
then
90-x = complement
180-x = supplement
.
90-x + 180-x = 110
270-2x = 110
-2x = -160
x = 80 degrees